OpenStudy (anonymous):

Quick physics question! Please help!! 3. A 0.200 kg air-track glider moving at 1.20 m/s bumps into a 0.600 kg glider at rest. a.) Find the total kinetic energy after collision if the collision is elastic. b.) Find the total kinetic energy after collision if the collision is completely inelastic.

OpenStudy (roadjester):

I explained elastic collisions to you before...

OpenStudy (anonymous):

sorry I forgot

OpenStudy (roadjester):

:( when you have an elastic collision, there are two conservation laws that you have to adhere to. Conservation of energy and conservation of momentum.

OpenStudy (roadjester):

\[\huge p_i=p_f\] where \[\huge p=mv\] so \[\huge m_1v_1=m_2v_2\] 1 represents the first glider and 2 represents the second glider which is at rest. This is the momentum part.

OpenStudy (anonymous):

ok..

OpenStudy (roadjester):

For an elastic collision, there has to be conservation of kinetic energy. So \[\huge {1 \over 2}m_1v_1^2={1 \over 2}m_2v_2^2\] same idea, 1 is the first glider, 2 is the second glider. there is NO PHYSICS involved. You said you were good at math. This isn't calculus. you know how to solve a system of equations right? well the first equation is the momentum, the second is kinetic energy both equations must be solved simultaneously

OpenStudy (roadjester):

have fun

OpenStudy (anonymous):

@ybarrap

OpenStudy (anonymous):

For a) I got 0.144 J because I used the kinetic energy formula. For b) well I'm kind of stuck on it

OpenStudy (anonymous):

@ybarrap ?

OpenStudy (ybarrap):

1/2 m v = 1/2 (m + M) v' pi = pf

OpenStudy (anonymous):

that would be for b?

OpenStudy (ybarrap):

yes

OpenStudy (anonymous):

ok so it would be 1/2(0.200)(1.20) = 1/2(0.200+0.600) * 0?

OpenStudy (ybarrap):

v' is not zero

OpenStudy (anonymous):

v is 1.20?

OpenStudy (ybarrap):

how did you get that?

OpenStudy (anonymous):

oh I'm not really sure what v is supposed to be just saw the 1.20 m/s and thought it was that

OpenStudy (ybarrap):